OpenStudy (anonymous):

The recursive formula for an arithmetic sequence is an = an-1 + d. If the common difference is -11 and a4 = 6, what is the value of a5? A. 17 B. -17 C. 5 D. -5

OpenStudy (anonymous):

\(a_{n}=a_{n-1}+d\\\large a_5=a_4+(-11)\)

OpenStudy (anonymous):

-5?????

OpenStudy (anonymous):

aye

OpenStudy (anonymous):

i was just about to say that just know so its 17

OpenStudy (anonymous):

i hold to use me calculator lol

OpenStudy (anonymous):

you was wrong

OpenStudy (anonymous):

...what was the answer?

OpenStudy (anonymous):

it wasnt a

OpenStudy (anonymous):

it doesnt tell me the answer i think it wwas b

OpenStudy (anonymous):

right?????

OpenStudy (anonymous):

I didn't say it was a. I said aye

OpenStudy (anonymous):

as in it was -5

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

omg lol its ok

OpenStudy (anonymous):

so yeah, for future reference aye=yes

OpenStudy (anonymous):

i thought you mean a instead of like aye

OpenStudy (freckles):

yes 6+(-11) isn't 17 it is 6+(-11)=-5 since -(11-6)=-(5)=-5

OpenStudy (anonymous):

ok